OpenStudy (anonymous):

PLEASE HELP I WANT OUT OF SUMMER SCHOOL put in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (anonymous):

y=mx+c Use that equation ^ Where m=slope And (5,-4)=(x,y)

OpenStudy (anonymous):

My program tells me to use Ax+By=C

OpenStudy (amistre64):

then use Ax + By = C

OpenStudy (amistre64):

slope tends to be: -a/b

OpenStudy (anonymous):

I have no idea how to do that! so that obviously didn't help

OpenStudy (amistre64):

then maybe you should stay in summer school :/

OpenStudy (amistre64):

your given the values of a and b in the form: -a/b pull them out and use them in the standard form your given the values of x and y in the form (x,y) use them in the standard form to determine the value of c

OpenStudy (anonymous):

sorry this is the first time ive been in public school. being homeschooled you dont learn much. so thanks

OpenStudy (amistre64):

lol, i learnt more teaching myself than anyone was ever able to teach me :)

OpenStudy (amistre64):

if we take the standard form ax + by = c and solve for y we get the slope intercept form y = -a/b x + c/b that -a/b is the slope of the equation; which they give you as -3/8

OpenStudy (amistre64):

therefore, a = 3 and b = 8 to fill in with

OpenStudy (amistre64):

3x + 8y = c ; use the x and y from the given point (5,-4) to determine c 5 -4 ------------ 15-32 = c the rest is clean up